# Attention Is Not What You Need: Revisiting Multi-Instance Learning for Whole Slide Image Classification

### v1

@@ 3.2 FocusMIL
Our main goal is to train a patch-level classifier from slide-level supervision while maintaining the standard multi-instance learning assumption and focusing on the causal factor $\bm{Z}_{c}$ that is invariant across slides. Formally, we assume
Assumption 1
@@
Corresponding to Assumption 1, $\bm{Z}_{s}$ varies between different patches, while $\bm{Z}_{B}$ varies between different slides.
Our proposed FocusMIL method encapsulates the two desiderata of WSI classification within weakly supervised variational inference, as shown in Figure 3. Firstly, raw input patches $\bm{p}$ are processed with pre-trained feature extractors into $\bm{x}$. Then, the encoder encodes $\bm{x}$ into low-dimensional latent variables $\bm{z}$, which is used for predicting the patch labels. The predicted patch labels are pooled with the max-pooling operation to comply with the standard MIL assumption. It is noteworthy that the max-pooling operation and the calculation of the loss function are conducted once per slide, and the encoder and classifier are optimized for the total loss of multiple bags within the mini-batch. This is fundamentally different from most attention-based WSI algorithms, where the optimization is restricted to one slide per epoch.
The Importance of the Standard MIL Assumption
@@
Models that respect the MIL assumptions predict bag labels solely based on the positive concept, e.g., the presence of tumour cells  (Foulds and Frank 2010; Dietterich, Lathrop, and Lozano-Pérez 1997). The two existing models that respect the MIL assumptions, mi-Net and CausalMIL, both use max pooling that aligns with the MIL problem setup. We argue that max-pooling effectively focuses the model on the causal factor $\bm{Z}_{c}$ for predicting the instance label. Due to the constraints of max-pooling, the model must base the bag label prediction solely on the most positive instance in the bag. If the classifier were to use $\bm{Z}_{B}$ to determine the instance label, because the bag-inherited information, such as staining conditions, is not entirely consistent across positive slides and differs on negative slides, the model would completely fail to fit the training set completely. Therefore, models that follow the MIL assumption must learn features such as the morphological structure of tumour cells, $\bm{Z}_{c}$, to fit the training set. A significant advantage is that models respecting MIL assumptions fitted well on the training set will achieve better patch-level classification results.
@@
For WSI classification, the morphology of tumour cells may be different across slides due to the varying malignancy levels of cancer in patients (Zhang et al. 2023; Qu et al. 2024; Jögi et al. 2012). As shown in Figure 2, some slides contain large tumour areas with significant morphological differences from normal cells. In contrast, other slides have smaller tumour areas where the tumour cells have higher differentiation and less pronounced differences from normal cells (Qu et al. 2024), which are referred to as “hard instances” in some literature (Tang et al. 2023). When multiple bags are in the mini-batch, the encoder and classifier will forward propagate multiple bags simultaneously. To ensure the model can accurately predict slides with different degrees of cancerous malignancy within the batch, the optimization process must consider the features of both salient and hard instances, thereby learning a better classification boundary. Consider the following scenario: if the model established a classification boundary based only on salient instances, it would not be able to accurately classify slides of lower malignancy, resulting in a high loss.
Batch size $>1$ also contributes to the disentanglement of $\bm{Z}_{c}$ and $\bm{Z}_{B}$. Staining conditions vary across slides. When faced with tumour cells of different colours (e.g., tumour cell patch stained purple and tumour cell patch stained pink), the model can only rely on $\bm{Z}_{c}$ for prediction. If it depends on the staining information $\bm{Z}_{B}$ inherited from the slide, it will not be able to achieve accurate predictions (Zhang et al. 2022b).
Forward Armotized Variational Inference
Max-pooling and batch optimization alone are insufficient for WSI classification. This is evident from the performance of a batched version of the classical mi-Net (Wang et al. 2018). mi-Net often correctly classifies the bag but fails to classify its patches during training. Furthermore, both bag and patch classification frequently fail entirely on the testing set.
We argue that this is because mi-Net, due to the flexibility of neural networks, may learn extremely complex decision surfaces that overfit the training data (Raff and Holt 2023) and fail to effectively learn latent causal representations.
We employ a variational inference with forward KL divergence $\bm{z}$ to address this issue. Specifically,
$D_{KL}[p(\bm{x},\bm{z})||q(\bm{x},\bm{z})]=$ $\mathbb{E}_{p(\bm{x},\bm{z})}[\log\frac{p(\bm{x},\bm{z})}{k(\bm{x})}]$
$-\mathbb{E}_{p(\bm{x},\bm{z})}[\log q(\bm{z}|\bm{x})],$ (3)
where $k(\bm{x})$ is a resampled distribution of the patches (Ambrogioni et al. 2019). As only the second term depends on the variational distribution, optimizing the forward KL divergence is equivalent to optimizing
$\mathcal{L}_{KL}=-\mathbb{E}_{p(\bm{x},\bm{z})}[\log q(\bm{z}|\bm{x})].$ (4)
The computation of the gradient in Equation 4 is straightforward as the expectation is defined for $p$ while the gradient is taken with respect to $q$.
For simplicity, we set the prior distribution as $p(\bm{z})=\mathcal{N}(0,\bm{I})$ and restrict the posterior as a factorized Gaussian encoder. Finally, the optimization objective can be obtained as:
$\mathcal{L}=$ $\ \log p(Y\mid\boldsymbol{z}^{*})-\mathbb{E}_{p(\bm{x},\bm{z})}\left[\log q(\boldsymbol{z}^{*}\mid\boldsymbol{x}^{*})\right]$
$\ -\beta\cdot\mathbb{E}_{q(\bm{z}|\bm{x})}\left[\log\frac{q(\boldsymbol{z}^{*}\mid\boldsymbol{x}^{*})}{p(\boldsymbol{z}^{*})}\right],$ (5)
where $\beta$ is a hyperparameter for balancing the classifier and variational loss, $\boldsymbol{z}^{*}$ corresponds to the most positive latent representation of a bag obtained with max-pooling with respect to $p(Y|\bm{z}^{*})$, and $\boldsymbol{x}^{*}$ corresponds to the original feature of $\bm{z}^{*}$. A noteworthy fundamental difference from CausalMIL (Zhang et al. 2022b) is that the utilization of forward KL entirely eliminates the need to optimize a decoder $p(\bm{x}|\bm{z})$, which is unnecessary for WSI classification.
@@ 4 Experiments
4.1 Datasets, Metrics, and Baselines
Camelyon16 dataset is widely used for metastasis detection in breast cancer (Bejnordi et al. 2017). The dataset consists of 270 training and 129 testing WSIs, which yield roughly 2.7 million patches at 10× magnification (the second level in multi-resolution pyramid). Pixel-level labels are available for tumour slides. Every WSI is cropped into $256\times 256$ patches without overlap, and background patches are discarded. A patch is labeled positive if it contains $25\%$ or more cancer areas. The numbers of tumorous versus normal patches are imbalanced as the positive patch ratios of positive slide in the training and testing sets of Camelyon16 are approximately 8.8% and 12.7%, respectively.
TCGA-NSCLC includes two subtypes of lung cancer: Lung Adenocarcinoma (LUAD) and Lung Squamous Cell Carcinoma (LUSC), with a total of 1,054 diagnostic digital slides. Only slide-level labels are available for this dataset. We directly used the patches released by Li et al. (Li, Li, and Eliceiri 2021).
Semi-Synthetic Dataset
To test whether existing WSI classification methods respect the standard MIL assumption, we propose the Camelyon16 Standard-MIL test dataset, inspired by the ideas in (Raff and Holt 2023). Specifically, in the training set, we introduce poison by randomly selecting $20\%$ of the patches in the normal slides and increasing the intensity of their green channel. In the test set, we randomly select $20\%$ of the patches in tumour slides to introduce poison in the same way. A MIL model cannot legally learn to use the poison signal because it occurs only in normal slides (Raff and Holt 2023). If a model has a training AUC $>0.5$, but a test AUC $<0.5$, it relies on poison for prediction and does not respect the MIL assumption.
Evaluation Metrics
Due to the high level of class imbalance at the patch level, we report Area Under the Precision-Recall Curve (AUCPR) and F1 score for evaluation. For slide-level classification, we report AUC and Accuracy as the tumour and normal classes are balanced.
Baselines
We compare our method with recently published baselines, including ABMIL (Ilse, Tomczak, and Welling 2018), DSMIL (Li, Li, and Eliceiri 2021), TransMIL (Shao et al. 2021), DFTD-MIL (Zhang et al. 2022a), IBMIL (Lin et al. 2023), and CausalMIL (Zhang et al. 2022b). The first five methods are based on attention mechanisms, while CausalMIL is an instance-based method which has not been previously applied to WSI classification, to the best of our knowledge. For the Camelyon16, TCGA-NSCLC, and Camelyon16 Standard-MIL test datasets, we used a ResNet-18 (He et al. 2016) pretrained on ImageNet (Deng et al. 2009) to extract features. In addition, we also use features extracted by the powerful CTransPath (Wang et al. 2022) for the Camelyon16 dataset to conduct a more comprehensive comparison of various MIL methods.
4.2 Implementation details
The encoder of FocusMIL consists of a neural network with one hidden layer, with ReLU as the activation function. We use AdamW optimizer with an initial learning rate of $1\times 10^{-4}$ to update the model weights during the training. The dimension of the latent factor is set to 35. The mini-batch size for training FocusMIL model is 3.
For slide-level prediction, we apply max-pooling and mean-pooling to aggregate instance prediction scores for the Camelyon16 and TCGA-NSCLC datasets, respectively.
For the attention-based methods, we built other models based on their officially released codes and conducted grid searches for key hyperparameters. For CausalMIL (Zhang et al. 2022b), both its encoder and decoder are set up as neural networks with hidden layer neurons 128 and two other fully-connected layers are used to carve factorized prior distribution conditioned on the bag information. All the experiments are conducted with a Nvidia RTX4090 GPU. For more details, please refer to our code in the supplementary.
Table 1: Performance Comparison of MIL Methods using Resnet-18 and CTransPath Features. The subscripts are the corresponding $95\%$ confidence intervals. The most superior performance is highlighted in bold, while the second-best performance is indicated by underlining.
@@ 4.2 Implementation details
ResNet-18 ImageNet pretrained ABMIL 0.8052 (CI: 0.7492, 0.8612) 0.8152 (CI: 0.7743, 0.8560) 0.1449 (CI: 0.0449, 0.2449) 0.1570 (CI: 0.0658, 0.2482)
DSMIL 0.7733 (CI: 0.7167, 0.8300) 0.7984 (CI: 0.7653, 0.8316) 0.2693 (CI: 0.0409, 0.4976) 0.2468 (CI: 0.1147, 0.3790)
TransMIL 0.8352 (CI: 0.7807, 0.8898) 0.8127 (CI: 0.7865, 0.8390) - -
DTFD-MIL 0.8619 (CI: 0.8530, 0.8709) 0.8032 (CI: 0.7497, 0.8567) 0.2079 (CI: 0.1592, 0.2566) 0.1640 (CI: 0.1388, 0.1891)
IBMIL 0.8442 (CI: 0.8329, 0.8555) 0.7906 (CI: 0.7465, 0.8347) 0.2587 (CI: 0.2138, 0.3037) 0.1948 (CI: 0.1499, 0.2398)
CausalMIL 0.8293 (CI: 0.8147, 0.8439) 0.8328 (CI: 0.8275, 0.8381) 0.4276 (CI: 0.4016, 0.4537) 0.4025 (CI: 0.3855, 0.4197)
Ours 0.8596 (CI: 0.8325, 0.8867) 0.8375 (CI: 0.8154, 0.8597) 0.3971 (CI: 0.3550, 0.4392) 0.4109 (CI: 0.3621, 0.4598)
Ctranspath SRCL ABMIL 0.9659 (CI: 0.9624, 0.9694) 0.9375 (CI: 0.9306, 0.9444) 0.3884 (CI: 0.3335, 0.4433) 0.3402 (CI: 0.2497, 0.4306)
DSMIL 0.9299 (CI: 0.9048, 0.9551) 0.9094 (CI: 0.8931, 0.9256) 0.6639 (CI: 0.6587, 0.6690) 0.6086 (CI: 0.5986, 0.6186)
TransMIL 0.9603 (CI: 0.9357, 0.9849) 0.9281 (CI: 0.8935, 0.9627) - -
DTFD-MIL 0.9739 (CI: 0.9722, 0.9756) 0.9515 (CI: 0.9472, 0.9559) 0.4535 (CI: 0.4315, 0.4756) 0.3350 (CI: 0.3072, 0.3628)
IBMIL 0.9716 (CI: 0.9691, 0.9740) 0.9562 (CI: 0.9476, 0.9649) 0.4285 (CI: 0.3977, 0.4594) 0.2927 (CI: 0.2607, 0.3247)
CausalMIL 0.9720 (CI: 0.9700, 0.9741) 0.9515 (CI: 0.9434, 0.9596) 0.6634 (CI: 0.6506, 0.6763) 0.6710 (CI: 0.6637, 0.6782)
Ours 0.9731 (CI: 0.9725, 0.9738) 0.9609 (CI: 0.9609, 0.9609) 0.6902 (CI: 0.6821, 0.6983) 0.6893 (CI: 0.6772, 0.7014)
Figure 4: Comparision of slide and patch level classification performance on validation set during training

### v2

@@ 4.3 Revisiting Max-Pooling-Based MIL
Max-pooling-based MIL inherently aligns the standard multiple-example learning assumption (Dietterich et al., 1997; Raff & Holt, 2023). For WSI classification task, the main goal is to train a patch-level classifier from slide-level supervision while maintaining the standard multi-instance learning assumption and focusing on the causal factor $\bm{Z}_{c}$ that is invariant across slides. Formally, we assume
Assumption 4.1.
@@
Corresponding to Assumption 4.1, $\bm{Z}_{s}$ may vary among different patches, while $\bm{Z}_{B}$ may vary among slides.
The Importance of the Standard MIL Assumption.
@@
Models that respect the MIL assumptions predict bag labels solely based on the positive concept, e.g., the presence of tumour cells  (Foulds & Frank, 2010; Dietterich et al., 1997). We argue that max-pooling effectively focuses the model on the causal factor $\bm{Z}_{c}$ for predicting the instance label. Due to the constraints of max-pooling, the model must base the bag label prediction solely on the most positive instance in the bag. If the classifier were to use $\bm{Z}_{B}$ to determine the instance label, because the bag-inherited information, such as staining conditions, is not entirely consistent across positive slides and differs on negative slides, the model would completely fail to fit the training set. Therefore, models that follow the MIL assumption must learn features such as the morphological structure of tumour cells, $\bm{Z}_{c}$, to fit the training set. A significant advantage is that models respecting MIL assumptions will achieve better patch-level classification results and enhanced out-of-distribution generalization.
Why Max-Pooling-Based MIL Methods Sometimes Underperform?
We believe there are two main reasons why the current max-pooling-based MIL methods do not perform ideally in the WSI classification task. The first reason is that mainstream methods like mi-Net (Wang et al., 2018) tend to directly remember the instances in the bag to fit the training set when the positive concept is difficult to learn. As shown in Figure 3, mi-Net perfectly fits the training set at the slide-level, while it produces almost random predictions in the validation set and fails almost completely for the patch-level predictions. This suggests that the model overfits by memorizing instances within the bag. As mentioned in Section 4.1, the reason for the difficulty in learning the positive concept is usually that the tumour region in the slide is small and tumours cells in some slides resemble normal cells. Empirically, we found that this problem disappears when using features extracted by advanced foundation models such as CTransPath (Wang et al., 2022).
The second reason, as discussed in Section 4.1, is that the malignancy levels of tumour cells across different slides are different. Since existing max-pooling-based methods select the most positive instance within one bag in a batch, existing max-pooling-based methods may perform well on salient instances but poorly on hard instances. However, this problem is also a challenge for attention-based methods.
Figure 3: Training and validation metrics over epochs for the mi-Net model on the Camelyon16 dataset.
Figure 4: Overview of our FocusMIL with 2 slides in a mini-batch for illustration. Features are extracted from patches using a pre-trained feature extractor. The instance features are encoded into latent representations via a variational encoder. The classifier $P(Y|\bm{z})$ uses the inferred latent representations $\bm{z}$ to obtain patch-level prediction scores. Max-pooling is applied to the instance predictions of each MIL bag to derive the slide-level prediction.
5 FocusMIL
To address the two aforementioned challenges, we propose FocusMIL. Figure 4 illustrates the architecture of our model. Firstly, raw input patches $\bm{p}$ are processed with pre-trained feature extractors into $\bm{x}$. Then, the variational encoder encodes $\bm{x}$ into a multivariate Gaussian distribution with a prior of $\mathcal{N}(0,I)$. Using the reparameterization trick (Kingma & Welling, 2013), latent representations $\bm{z}$ are sampled from this distribution and utilized to predict patch-level labels. The patch-level predictions are aggregated with the max-pooling operation to comply with the standard MIL assumption. Another key difference between our method and mi-Net is that we optimize multiple bags simultaneously within a mini-batch. We will explain our method in the following sections.
Addressing Overfitting with Variational Encoder
Since max-pooling leads to one instance per bag being trained in one iteration, as discussed in Section 4.3, when the positive concept is difficult to learn, the model fits the training set by remembering the instance within the bag. To prevent the classifier from directly memorizing instances, we employ a variational encoder that encodes each instance into a distribution and resample from the distribution to obtain a latent representation. Inspired by Variational Autoencoders (VAEs) (Kingma & Welling, 2013), we use a KL loss to regularize the latent space, ensuring it remains smooth and meaningful. We set the prior distribution as $p(\bm{z})=\mathcal{N}(0,\bm{I})$.
Our final loss function is:
$\mathcal{L}=-\log p\left(Y\mid\bm{z}^{*}\right)+\beta\cdot\operatorname{KL}\left[q\left(\bm{z}\mid\bm{x}\right)\|p\left(\bm{z}\right)\right]$ (6)
where $\beta$ is a hyperparameter for balancing the classifier and KL loss, $\bm{z}^{*}$ corresponds to the most positive latent representation of a bag obtained with max-pooling with respect to $p(Y|\bm{z})$.
The key difference between our method and CausalMIL (Zhang et al., 2022b) is that our approach regularizes the latent space of all instances, whereas CausalMIL only regularizes the latent space of the most positive instance per bag. The regularization approach of CausalMIL results in a large number of instances’ latent spaces remaining unconstrained, making it hard to sample meaningful representations for those instances. Our subsequent experiments also demonstrate that CausalMIL cannot effectively mitigate the simple memorization of bags on more complex datasets. Additionally, since we believe that adhering to the standard MIL assumption inherently enforces the disentangling of $\bm{Z}_{c}$ and $\bm{Z}_{B}$, our method does not require instance reconstruction.
@@ 5 FocusMIL
For WSI classification, the morphology of tumour cells may be different across slides due to the varying malignancy levels of cancer in patients (Zhang et al., 2023; Qu et al., 2024; Jögi et al., 2012). When multiple ($n$) bags are in the mini-batch, the encoder and classifier will forward propagate multiple bags simultaneously. For each bag, max-pooling selects the most positive instance. Since the most positive patches in different tumour slides within a batch exhibit varying malignancy levels, the optimization process must consider the features of these $n$ patches, thereby learning a better classification boundary. Consider the following scenario: if the model established a classification boundary based only on salient instances, it would not be able to accurately classify slides of lower malignancy, resulting in a high loss.
Batch size $>1$ also contributes to the disentanglement of $\bm{Z}_{c}$ and $\bm{Z}_{B}$. Staining conditions vary across slides. When faced with tumour cells of different colours (e.g., tumour cell patch stained purple and tumour cell patch stained pink), the model can only rely on $\bm{Z}_{c}$ for prediction. If it depends on the staining information $\bm{Z}_{B}$ inherited from the slide, it will not be able to achieve accurate predictions (Zhang et al., 2022b).
@@ 6 Experiments
Table 1: Performance Comparison of Multiple MIL Methods using ResNet-18 and CTransPath Features. The subscripts are the corresponding $95\%$ confidence intervals. The best performance is in bold, while the second-best is underlined.
@@
ResNet-18 ABMIL 0.8052 (0.7492, 0.8612) 0.8152 (0.7743, 0.8560) 0.1449 (0.0449, 0.2449) 0.1570 (0.0658, 0.2482)
DSMIL 0.7733 (0.7167, 0.8300) 0.7984 (0.7653, 0.8316) 0.2693 (0.0409, 0.4976) 0.2468 (0.1147, 0.3790)
TransMIL 0.8352 (0.7807, 0.8898) 0.8127 (0.7865, 0.8390)
DTFD-MIL 0.8619 (0.8530, 0.8709) 0.8032 (0.7497, 0.8567) 0.2079 (0.1592, 0.2566) 0.1640 (0.1388, 0.1891)
IBMIL 0.8442 (0.8329, 0.8555) 0.7906 (0.7465, 0.8347) 0.2587 (0.2138, 0.3037) 0.1948 (0.1499, 0.2398)
mi-Net 0.8014 (0.7674, 0.8355) 0.8158 (0.7900, 0.8416) 0.4108 (0.3466, 0.4751) 0.3844 (0.3437, 0.4251)
CausalMIL 0.8092 (0.7730, 0.8454) 0.8281 (0.8128, 0.8435) 0.4845 (0.4653, 0.5036) 0.3650 (0.3469, 0.3823)
FocusMIL 0.8706 (0.8489, 0.8923) 0.8359 (0.8206, 0.8513) 0.4382 (0.3907, 0.4857) 0.3996 (0.3787, 0.4205)
CTransPath ABMIL 0.9659 (0.9624, 0.9694) 0.9375 (0.9306, 0.9444) 0.3884 (0.3335, 0.4433) 0.3402 (0.2497, 0.4306)
DSMIL 0.9299 (0.9048, 0.9551) 0.9094 (0.8931, 0.9256) 0.6639 (0.6587, 0.6690) 0.6086 (0.5986, 0.6186)
TransMIL 0.9603 (0.9357, 0.9849) 0.9281 (0.8935, 0.9627)
DTFD-MIL 0.9739 (0.9722, 0.9756) 0.9515 (0.9472, 0.9559) 0.4535 (0.4315, 0.4756) 0.3350 (0.3072, 0.3628)
IBMIL 0.9716 (0.9691, 0.9740) 0.9562 (0.9476, 0.9649) 0.4285 (0.3977, 0.4594) 0.2927 (0.2607, 0.3247)
mi-Net 0.9757 (0.9723, 0.9791) 0.9593 (0.9550, 0.9637) 0.6156 (0.6007, 0.6305) 0.5793 (0.5542, 0.6045)
CausalMIL 0.9700 (0.9661, 0.9740) 0.9531 (0.9412, 0.9650) 0.6640 (0.6509, 0.6770) 0.6107 (0.5762, 0.6452)
FocusMIL 0.9731 (0.9725, 0.9738) 0.9609 (0.9609, 0.9609) 0.6902 (0.6821, 0.6983) 0.6893 (0.6772, 0.7014)
Table 2: Localization FROC Performance Comparison. FocusMIL (1) and FocusMIL (3) refer to the FocusMIL models trained with a batch size of 1 and 3, respectively.
Method Localization FROC
ABMIL 0.2434
DTFD-MIL 0.3013
FocusMIL (1) 0.3462
FocusMIL (3) 0.3646
6.1 Datasets, Metrics, and Baselines
Camelyon16 dataset is widely used for metastasis detection in breast cancer (Bejnordi et al., 2017a). The dataset consists of 270 training and 129 testing WSIs.
Camelyon17 dataset consists of 1,000 WSIs from five hospitals. Due to the absence of labels in the test set, we only used the 500 slides from the training set. Following the setup of AEM (Zhang et al., 2024), we designated 200 slides from the fourth and fifth hospitals as the test set to evaluate the model’s OOD generalization performance. The remaining 300 slides were split into training and validation sets with an 8:2 ratio. For a more detailed discussion of the Camelyon datasets, as well as additional OOD experiment results, please refer to the Appendix.
Semi-Synthetic Dataset To test whether existing WSI classification methods respect the standard MIL assumption, we propose the Camelyon16 Standard-MIL test dataset, inspired by the ideas in (Raff & Holt, 2023). Specifically, in the training set, we introduce poison by randomly selecting $20\%$ of the patches in the normal slides and increasing the intensity of their green channel. In the test set, we randomly select $20\%$ of the patches in tumour slides to introduce poison in the same way. A MIL model cannot legally learn to use the poison signal because it occurs only in normal slides (Raff & Holt, 2023). If a model has a training AUC $>0.5$, but a test AUC $<0.5$, it relies on poison for prediction and does not respect the MIL assumption.
Evaluation Metrics Due to the high level of class imbalance at the patch level, we report Area Under the Precision-Recall Curve (AUCPR) and F1 score for evaluation. For slide-level classification on Camelyon16 dataset, we report AUC and Accuracy as the tumour and normal classes are balanced. We report AUC and F1-score on Camelyon17 dataset, as the classes are relatively imbalanced. On Camelyon16, we also evaluate tumour region localization performance by reporting free response operating characteristic curves (FROC) (Bejnordi et al., 2017a).
Baselines We compare our method with several recently published baselines, including ABMIL (Ilse et al., 2018), DSMIL (Li et al., 2021), TransMIL (Shao et al., 2021), DFTD-MIL (Zhang et al., 2022a), IBMIL (Lin et al., 2023), mi-Net (Wang et al., 2018), and CausalMIL (Zhang et al., 2022b). The first five methods use attention mechanisms, while mi-Net and CausalMIL are based on max-pooling. To the best of our knowledge, CausalMIL has not yet been applied for WSI classification. For the Camelyon16 dataset, we adopt ResNet18 (He et al., 2016) (pretrained on ImageNet (Deng et al., 2009)) and CTransPath (Wang et al., 2022) as feature extractors. For the Camelyon17 dataset, we directly utilize the features released by AEM (Zhang et al., 2024), which include features extracted by ResNet-18 and PathGen-CLIP (Sun et al., 2024). For the semi-synthetic dataset, ResNet-18 is used to extract features.
Table 3: Performance Comparison of MIL Methods on Camelyon17 Dataset using ResNet-18 and PathGen-CLIP Features.
Method Performance Slide-level
AUC F1-score
ResNet-18 ABMIL 0.7877 (0.7505, 0.8249) 0.7004 (0.6702, 0.7306)
DSMIL 0.7069 (0.5669, 0.8469) 0.6507 (0.5826, 0.7187)
TransMIL 0.6899 (0.6187, 0.7610) 0.6200 (0.5459, 0.6941)
DTFD-MIL 0.8192 (0.7964, 0.8419) 0.7242 (0.6960, 0.7524)
IBMIL 0.8221 (0.8059, 0.8383) 0.7233 (0.7022, 0.7444)
mi-Net 0.5066 (0.5022, 0.5110) 0.2901 (0.2419, 0.3383)
CausalMIL 0.5104 (0.4615, 0.5592) 0.4905 (0.3881, 0.5929)
FocusMIL 0.8719 (0.8619, 0.8818) 0.8335 (0.8273, 0.8398)
PathGen-CLIP ABMIL 0.9264 (0.9156, 0.9372) 0.8909 (0.8880, 0.8938)
DSMIL 0.9212 (0.9035, 0.9390) 0.8904 (0.8851, 0.8957)
TransMIL 0.9387 (0.9274, 0.9500) 0.8932 (0.8871, 0.8993)
DTFD-MIL 0.9447 (0.9341, 0.9553) 0.8870 (0.8788, 0.8952)
IBMIL 0.9410 (0.9308, 0.9512) 0.8854 (0.8764, 0.8944)
mi-Net 0.9442 (0.9307, 0.9577) 0.8882 (0.8819, 0.8946)
CausalMIL 0.9537 (0.9445, 0.9628) 0.8885 (0.8799, 0.8970)
FocusMIL 0.9529 (0.9456, 0.9602) 0.8974 (0.8875, 0.9073)